Briley, Inc., is expected to pay equal dividends at the end of each of the next two years. Thereafter, the dividend will grow at a constant annual rate of 4.6 percent, forever. The current stock price is $51.

  

What is next year’s dividend payment if the required rate of return is 13 percent? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  

  Dividend payment
Answer #1

Next Year's dividend payment will be $ 4.47.

As per the dividend discount model, current stock price is the present value of future dividends.
Step-1:Present value of equal dividend of next 2 years
Present value = D*(1+0.13)^-1 + D*(1+0.13)^-2
= 0.884956 D + 0.783147 D
= 1.668102 D
Step-2:Present value of dividends after year 2
Present value = D2*(1+g)/(Ke-g)*DF2 Where,
= D*(1+0.046)/(0.13-0.046)*0.783147 D2 = Dividend of year 2 = D
= 9.752041 D g = Growth rate = 4.60%
Ke = Required return = 13%
DF2 = Discount factor = 0.783147
Step-3:Calculation of next year's dividend
As per dividend discount model, price of stock is the sum of present value of dividends.
As per question,
Current stock price = Sum of present value of future dividends
$    51.00 = 1.668102 D + 9.752041 D
$    51.00 = 11.42014 D
D = $       4.47
So, next years dividend will be $ 4.47